OpenStudy (anonymous):

the product of 2 consecutive odd integers is 143. find the integers?

Directrix (directrix):

Let x be an odd integer. Then, the next consecutive odd integer is x + 2.

Directrix (directrix):

The product of the two integers is the following: x ( x + 2) = 143

Directrix (directrix):

@Garrett1234 Would you multiply this, please? x ( x + 2) = 143

Directrix (directrix):

Use the distributive property to multiply x by (x + 2)

OpenStudy (anonymous):

x^2 + 2x = 143 right? @Directrix

Directrix (directrix):

Yes.

OpenStudy (anonymous):

Yes. Now bring the 143 over to the other side and solve the quadratic.

Directrix (directrix):

Now, it is time to factor x^2 +2x - 143 = 0.

Directrix (directrix):

Look for two numbers that multiply to -143 and add to +2.

Directrix (directrix):

@Garrett1234 What is ? below. (x + 13) ( x - ? ) = 0

OpenStudy (skullpatrol):

@Directrix How did you pull 13 "out of a hat"?

OpenStudy (anonymous):

(x+13) (x-11) = 0 correct? @Directrix

Directrix (directrix):

@Garrett1234 Yes. So far, so good.

Directrix (directrix):

It is Zero Product Property time now. If the product of two quantities is zero, then one or the other or both (the factors) = 0. (x+13) (x-11) = 0 means x + 13 = 0 OR x - 11 = 0 @Garrett1234 Would you solve each of those two equations for x?

OpenStudy (anonymous):

\[(2 n+1) (2 n+3)=143 \]\[\{n\to -7,n\to 5\} \]{-13,-11} and {11,13}

Directrix (directrix):

@Garrett1234 Here's where left off: (x+13) (x-11) = 0 means x + 13 = 0 OR x - 11 = 0 @Garrett1234 Would you solve each of those two equations for x?

Directrix (directrix):

@skullpatrol @robtobey

OpenStudy (anonymous):

So, x=-13, or 11 right? @Directrix

Directrix (directrix):

Correct. x + 13 = 0 OR x - 11 = 0 x = -13 OR x = 11 @Garrett1234

Directrix (directrix):

@Garrett1234 What is ? below? If x = -13, then the next consecutive odd integer is -13 + 2 = ?

OpenStudy (anonymous):

-11 right? @Directrix

Directrix (directrix):

Correct, @Garrett1234 So, one answer pair is the integers: -13 and -11 But, there is another.

Directrix (directrix):

@Garrett1234 What is ? below. If x = 11, then the next consecutive odd integer is 11 + 2 = ?

OpenStudy (anonymous):

So the integers could either be -13, -11 or 11, 13 correct? @Directrix

Directrix (directrix):

@Garrett1234 Yes, but remember that they come in pairs. Avoid running them all together as 4 numbers. The complete answer must have both pairs of integers: -13 and -11 --> one pair AND 11 and 13 --> second pair

Directrix (directrix):

@Garrett1234 Questions?

OpenStudy (anonymous):

No, you have been tons of help, thank you so much! You explained it all very clearly and took it step by step. @Directrix Although i do have one more problem that seems a little more advanced.

Directrix (directrix):

Glad to help. Post your next question in a new thread. We seem to have lots of Answerers interested in this work. That's great. I will follow you to the new thread when you have the next question posted. Okay? @Garrett1234

OpenStudy (anonymous):

Okay, and once again thanks! @Directrix

Directrix (directrix):

You are welcome. I just saw that this is your first posted question in the mathematics section. And, a BIG WELCOME to OPEN STUDY MATH. If you have any questions about the site, just hit the @Directrix wherever you are working here. If I am online, I will find you and help. Or, you can send a private message, and I will try my best to help out. @Garrett1234

Directrix (directrix):

@aayush12345